[Ian, Torrey, Sander]
We moved the sleds for the rubidium vapor cell locking experiment as well as the sled that holes the SHG optics for the 1560 nm to 780 nm conversion to the north most small table in the atoms lab (B11A). We disconnected all of the cables in B102 and moved the sleds as well as the 780 nm laser (McCuller Lab SN: S2400003) and the laser current and temp controllers. We also moved the temp controller for the vapor cell. I haven't moved any of the cables for any of the instruments of hooked anything up yet. I plan on trying to do that sometime on Friday. Then I will briefly turn on the laser to make sure it works.
Because we moved the laser we need to update the standard operating procedures (SOP) for the laser so that it refers to the new lab. We also need to update the Caltech property tag location. The Caltech property tag is (C0E000083455). I have already updated the wiki for the location. As well as added a link to the SOP there.